OpenStudy (anonymous):

Rusty has a boat goes 12 mi. an hour. Find the current speed of the river if he can go 3 mi. upstream in the same time as he goes 5 mi. downstream. Please show work

OpenStudy (anonymous):

your speed*time=distance traveled so: 3=(12-x)t and 5=(12+x)t solve for t in both equations 3/(12-x)=t 5/(12+x)=t substitute: 3/(12-x)=5/(12+x) cross multiply 3(12+x)=5(12-x) 36+3x=60-5x 8x=24 x=3 so stream current is 3 miles/hr downstream

OpenStudy (anonymous):

thats what I got. I just wanted a second opinion!!! Thanks!!!

OpenStudy (anonymous):

no prob